It takes a cabin cruiser 3 hours longer to travel 50kms than it takes a power boat to travel 60kms. The cruiser travels one third the speed of the power boat. Find the speed of each boat.
Answer by stanbon(48558)

It takes a cabin cruiser 3 hours longer to travel 50kms than it takes a power boat to travel 60kms. The cruiser travels one third the speed of the power boat. Find the speed of each boat.
-------------
Cruiser DATA:
distance = 50 km ; rate = x kph ; time = 50/x hrs
--------------
Power Boat DATA:
distance = 60 km ; rate = 3x kph ; time = 60/(3x) hrs
-----------------
EQUATION:
cruise time = power boat time + 3 hrs.
50/x = 60/3x + 3
--------
Multiply thru by 3x to get:
150 = 60 + 9x
210 = 9x
x = 70/3 = 23 1/3 kph (cruiser speed)
3x = 70 kph (power boat speed)
==================
Cheers,
Stan H.